Question 59 the equivalent units of production as per the data given; Units started 15,000 Units completed 10,000 Units in process 5,000 (60 % completed) Question 60 Calculate the equivalent units of production as per the data given; Units started 5,000 Units completed -4,000 Units in process-1,000 (20 % completed)
Answer #1

Question 59

Equivalent units of production = units completed + units in process = 4000 +1000 =5000

Question 60

Equivalent units of production = 4000+1000 = 5000